# First-order definability of Campana Points and Darmon Points in algebraic function fields in one variable over number fields

### v1

@@ Abstract.
We give first-order definitions of Campana and Darmon points in algebraic function fields in one variable over number fields. These sets are geometric generalizations of $n$-full integers (integers whose nonzero valuations are at least $n$) and perfect $n$th powers, respectively, to more general algebraic varieties. For this we exploit the theory of quadratic Pfister forms, which were used in [2] and [4] to extend to the case of algebraic function fields in one variable the methods used by Koenigsmann in [17]. These methods had already been generalized to arbitrary global fields by Park in [28] and Eisenträger & Morrison in [10], and the author had already exploited these methods to find first-order definitions of Campana points (in [33]) and Darmon points (in [34], with Handley) in the context of number fields. With the newly expanded version of these methods, we now transfer those results to the new context of algebraic function fields in one variable over number fields.
@@ 1. Introduction
If $R$ is an integral domain, $n\in\mathbb{Z}_{\geq 1}$ and $A\subseteq R^{n}$, we say that $A$ is diophantine (or existentially definable) over $R$ if there exists $m\in\mathbb{Z}_{\geq 0}$ and a polynomial $P\in R\left[x_{1},\cdots,x_{m},y_{1},\cdots,y_{n}\right]$ such that $A=\left\{\vec{y}\in R^{n}:\exists x_{1},\cdots,x_{m}\in R\left(P\left(x_{1},\cdots,x_{m},\vec{y}\right)=0\right)\right\}$. The understanding of diophantine sets is intimately related to the question about the decidability of the positive existential theory of the ring in question. Hilbert’s Tenth Problem originally asked this question when $R=\mathbb{Z}$, which was negatively answered by Matiyasevich in [21], building its argument from Davis, Putnam, and Robinson’s research on exponential diophantine equations in [5]. Even though for various other rings an answer has also been attained (see for example [6], [14], [22], [19]), the decidability of the positive existential theory of $\mathbb{Q}$ and, more generally, any global field, remains unanswered. Given an integral domain $R$, the question about the decidability of the positive existential theory of $R$ is commonly known as Hilbert’s Tenth Problem over $R$.
If $R^{\prime}$ is a subring of $R$, the diophantineness of $R^{\prime}$ over $R$ allows to transfer a negative answer to Hilbert’s Tenth Problem over $R^{\prime}$ to $R$ (see for example [16, Proposition 2.1]), which explains the deep interest in the study of diophantine sets. In an attempt to contribute towards an existential definition of $\mathbb{Z}$ in $\mathbb{Q}$, Koenigsmann showed in [17] that $\mathbb{Z}$ is universally defined in $\mathbb{Q}$ (i.e. it is the complement of a diophantine set), by developing methods and results that were later generalized to prove analogous statements on more arbitrary global fields (see for example [28], [10], [7], [25], [3]). Since the question about the undecidability of the positive existential theory of $\mathbb{Q}$ (or, more generally, a global field $K$) seems currently out of reach, research has tried an “intermediate step” approach, by studying intermediate subrings (some instances being [31], [11], [29], [8], [9]). Motivated by this approach, the author offered a $\forall\exists$-definition for every member of a sequence of sets that interpolates between a number field and its set of $S$-integers. These sets are known as the Campana points (see [33]). Together with Handley, the author was able to extend his methods to give a similar approach for Darmon points in [34].
Recently Becher, Daans, and Dittmann were able to further generalize Koenigsmann’s methods to the case of algebraic function fields in one variable (see [2] and [4]). This motivated the author to transfer his previous results to this new scenario, and in this paper we prove:
@@
In the upcoming section we will see the specific definitions for Campana and Darmon points. It will be important to note from the beginning thay these sets are given in terms of a finite set of $\mathbb{Z}$-valuations that are trivial on the base field. To be more specific, given an algebraic function field $F$ in one variable over a number field $K$, given $n\in\mathbb{Z}_{\geq 1}$ and given a finite set $S$ of $\mathbb{Z}$-valuations on $F$ that are trivial on $K$, we will define the sets $C_{F/K,S,n}$ (the set of $n$-Campana points of $F/K$ with respect to $S$) and $D_{F/K,S,n}$ (the set of $n$-Darmon points of $F/K$ with respect to $S$). As we will see, the first-order descriptions we will give will be uniform with respect to all possible choices of $S$ (cf. Remark 8.11 and Remark 8.13).
The above remark is important because, by [4, Corollary 5.9], if $K$ is a number field and $F$ is an algebraic function field in one variable over $K$, every computably enumerable relation on $F$ is $\exists\forall$-definable. In particular, since every decidable subset is both computably enumerable and has a computably enumerable complement, then all decidable subsets are both $\forall\exists$-definable and $\exists\forall$-definable. Since Campana points and Darmon points are decidable, we can get our first-order characterization a priori. The difference here is that our explicit approach towards a specific construction of their defining formulas will reveal to be uniform with respect to all finite subsets $S$ of $\mathbb{Z}$-valuations trivial on the base field in terms of which these sets are defined. More specifically, the sequence of Proposition 8.2, Proposition 8.3, and Proposition 8.6 will allow for a parametrization of all finite subsets of such places in a way that will allow for a uniform control on which of those sets $S$ is the one involved in the defining formula.
@@ 1.1. Understanding Campana and Darmon points
  • Darmon, if for all $\alpha\in\mathcal{A}$ with $\varepsilon_{\alpha}=1$ we have $n_{v}\left(\mathcal{D}_{\alpha},P\right)=0$, and for all $\alpha\in\mathcal{A}$ with $\varepsilon_{\alpha}<1$ and $n_{v}\left(\mathcal{D}_{\alpha},P\right)>0$ we have $n_{\alpha}\mid n_{v}\left(\mathcal{D}_{\alpha},P\right)$.
For example, as shown in [34, §3.1], when $X=\mathbb{P}^{1}_{\mathbb{Q}}$, $\mathcal{X}=\mathbb{P}^{1}_{\mathbb{Z}}$, $D\coloneqq\left(1-\frac{1}{n}\right)\left\{x_{0}=0\right\}+\left\{x_{1}=0\right\}$ for some $n\in\mathbb{Z}_{\geq 1}$, and $S=\left\{\infty\right\}$, then the corresponding Darmon points in the affine line are the integral perfect $n$th powers up to sign. If instead $D\coloneqq\left(1-\frac{1}{n}\right)\left\{x_{0}=0\right\}+\left(1-\frac{1}{n}\right)\left\{x_{1}=0\right\}$ then we get rational $n$th powers up to sign.
If $K$ is any number field, $S$ is a finite set of places of $K$ containing the archimedean ones, $X=\mathbb{P}^{1}_{K}$, $\mathcal{X}=\mathbb{P}^{1}_{\mathcal{O}_{K,S}}$, $n\in\mathbb{Z}_{\geq 1}$ and $D=\left(1-\frac{1}{n}\right)\left\{x_{1}=0\right\}$, then calling $C_{K,S,n}$ and $D_{K,S,n}$ the corresponding Campana and Darmon points, respectively, as shown in [33, §2.2] and [34, §3.2] we get
@@
These are the sets that were explored in [33] and [34] and will be generalized in our new context.
@@ 1.1.1. Other contexts in which these sets have been studied
Even though these sets were mainly motivated by their interest in proving Manin-type conjectures and also in counting problems with height restrictions (see for instance [1], [24], [30], [37]), they have also been studied in the context of definability and decidability. For example, in [15] Garcia-Fritz, Pasten, and Pheidas showed that Kollár’s conjecture is incompatible with the diophantineness of Campana points in $\mathbb{C}\left(z\right)$, except for the trivial cases. We state the conjecture below. For each $n\in\mathbb{Z}_{\geq 1}$ we let $\mathbb{C}\left[z\right]_{n}$ be the $\left(n+1\right)$-dimensional $\mathbb{C}$-vector subspace of $\mathbb{C}\left[z\right]$ spanned by $\left\{1,x,\cdots,x^{n}\right\}$. We endow this set with the Zariski topology of $\mathbb{C}^{n+1}$.
@@
If $D$ is a diophantine subset of $\mathbb{C}\left(z\right)$ containing a Zariski open subset of $\mathbb{C}\left[z\right]_{n}$ for infinitely many $n\in\mathbb{Z}_{\geq 1}$, then $\mathbb{C}\left(z\right)\setminus D$ is finite.
@@
$C_{S,n}$ $\coloneqq\left\{f\in\mathbb{C}\left(z\right):\forall\alpha\in\mathbb{P}^{1}_{\mathbb{C}}\left(\mathbb{C}\right)\setminus S\,\left(\nu_{\alpha}\left(f\right)\in\mathbb{Z}_{\geq 0}\cup\mathbb{Z}_{\leq-n}\right)\right\},$
$D_{S,n}$ $\coloneqq\left\{f\in\mathbb{C}\left(z\right):\forall\alpha\in\mathbb{P}^{1}_{\mathbb{C}}\left(\mathbb{C}\right)\setminus S\,\left(\nu_{\alpha}\left(f\right)\in\mathbb{Z}_{\geq 0}\cup n\mathbb{Z}\right)\right\},$
@@
It is clear that $C_{S,1}=D_{S,1}=\mathbb{C}\left(z\right)$ is diophantine in $\mathbb{C}\left(z\right)$. Moreover, $C_{\emptyset,\infty}=D_{\emptyset,\infty}=\mathbb{C}$ is also diophantine in $\mathbb{C}\left(z\right)$ because, in general, if $L$ is a large field and $F$ is an algebraic function field in one variable over $L$, then $L$ is diophantine over $F$, by [18, Theorem 2]. We recall that a field $L$ is large if and only if every smooth curve over $L$ has either infinitely many or none $L$-rational points.
In [15, Proposition 1.13] Garcia-Fritz, Pasten, and Pheidas showed that, in the case of Campana points, these are the only instances of diophantineness that can occur if Conjecture 1.2 is true. For the sake of completeness, let us now show how this argument can be adapted to get an analogous proposition for Darmon points.
@@ 3. $\mathbb{Z}$-valuations on algebraic function fields in one variable
Recall that a $\mathbb{Z}$-valuation $v$ on a field $K$ is a discrete valuation such that for every $n\in\mathbb{Z}$ there exists $x\in K$ such that $n=v\left(x\right)$. If $L$ is a finite field extension of $K$ and $w$ is a $\mathbb{Z}$-valuation on $K$, we say that $w$ lies over $v$ (or that $v$ lies under $w$) if and only if $\mathcal{O}_{w}\cap K=\mathcal{O}_{v}$. In this case, there exists $e\left(w\mid v\right)\in\mathbb{Z}_{\geq 1}$ such that $w\left(x\right)=e\left(w\mid v\right)v\left(x\right)$ for every $x\in K$. The number $e\left(w\mid v\right)$ is called the ramification index of $w$ over $v$. We say that $v$ is unramified in $L$ if $e\left(w\mid v\right)=1$ for every $\mathbb{Z}$-valuation on $L$ lying over $K$. It is well-known that any $\mathbb{Z}$-valuation on $K$ lies under only finitely many (and at least one) $\mathbb{Z}$-valuations on $L$.
We refer the reader to [26] and [13] for more properties of valuations. In our case, we will be considering $\mathbb{Z}$-valuations on algebraic function fields in one variable.
@@
$C_{F/K,S,n}$ $\coloneqq\left\{x\in F:v\left(x\right)\in\mathbb{Z}_{\geq 0}\cup\mathbb{Z}_{\leq-n}\text{ for all $v\in\mathcal{V}\left(F/K\right)\setminus S$}\right\},$
$D_{F/K,S,n}$ $\coloneqq\left\{x\in F:v\left(x\right)\in\mathbb{Z}_{\geq 0}\cup n\mathbb{Z}\text{ for all $v\in\mathcal{V}\left(F/K\right)\setminus S$}\right\}.$
@@ 4. Quadratic Forms and Witt Equivalence
Assume fist $\operatorname{Char}\left(K\right)\neq 2$. By the diagonalization theorem for quadratic forms on characteristic different from $2$ ([12, Proposition 7.29]) and the fact that $q$ is non-singular, $q$ is isometric to the orthogonal product of quadratic forms $K\to K$ given by $x\mapsto ax^{2}$ for every $x\in K$, where $a\in K^{\times}$. Calling $\left\langle a\right\rangle$ such a quadratic form for a given $a\in K^{\times}$, the first conclusion follows from the fact that $\left\langle a\right\rangle\perp\left\langle-a\right\rangle$ is hyperbolic ([12, Example 7.26]). The second conclusion is a direct consequence of hyperbolic forms always being even-dimensional.
The proof when $\operatorname{Char}\left(K\right)=2$ is essentially the same; with the additional task of dealing with the more complicated version of the orthogonal factorization of quadratic forms on fields of characteristic $2$ ([12, Proposition 7.31]), since we have to also consider orthogonal factors $K\times K\to K$ of the form $\left(x,y\right)\mapsto ax^{2}+xy+by^{2}$ with $a,b\in K$. Calling $\left[a,b\right]$ such a quadratic form for given $a,b\in K$, [12, Example 7.23] gives $\left[a,b\right]\perp\left[a,-b\right]\cong\left[a,0\right]\perp\mathbb{H}\left(K\right)$, and [12, Example 7.5] gives $\left[a,0\right]\cong\mathbb{H}\left(K\right)$. $\blacksquare$

### v2

@@ Abstract.
We give first-order definitions of Campana and Darmon points in algebraic function fields in one variable over number fields. These sets are geometric generalizations of $n$-full integers (integers whose nonzero valuations are at least $n$) and perfect $n$th powers, respectively, to more general algebraic varieties. For this we exploit the theory of quadratic Pfister forms, which were used in [undefa] and [undefc] to extend to the case of algebraic function fields in one variable the methods used by Koenigsmann in [undefp]. These methods had already been generalized to arbitrary global fields by Park in [undefaa] and Eisenträger & Morrison in [undefi], and the author had already exploited these methods to find first-order definitions of Campana points (in [undefaf]) and Darmon points (in [undefag], with Handley) in the context of number fields. With the newly expanded version of these methods, we now transfer those results to the new context of algebraic function fields in one variable over number fields.
@@ 1. Introduction
If $R$ is an integral domain, $n\in\mathbb{Z}_{\geq 1}$ and $A\subseteq R^{n}$, we say that $A$ is diophantine (or existentially definable) over $R$ if there exists $m\in\mathbb{Z}_{\geq 0}$ and a polynomial $P\in R\left[x_{1},\cdots,x_{m},y_{1},\cdots,y_{n}\right]$ such that $A=\left\{\vec{y}\in R^{n}:\exists x_{1},\cdots,x_{m}\in R\left(P\left(x_{1},\cdots,x_{m},\vec{y}\right)=0\right)\right\}$. The understanding of diophantine sets is intimately related to the question about the decidability of the positive existential theory of the ring in question. Hilbert’s Tenth Problem originally asked this question when $R=\mathbb{Z}$, which was negatively answered by Matiyasevich in [undeft], building its argument from Davis, Putnam, and Robinson’s research on exponential diophantine equations in [undefd]. Even though for various other rings an answer has also been attained (see for example [undefe], [undefm], [undefu], [undefr]), the decidability of the positive existential theory of $\mathbb{Q}$ and, more generally, any global field, remains unanswered. Given an integral domain $R$, the question about the decidability of the positive existential theory of $R$ is commonly known as Hilbert’s Tenth Problem over $R$.
If $R^{\prime}$ is a subring of $R$, the diophantineness of $R^{\prime}$ over $R$ allows to transfer a negative answer to Hilbert’s Tenth Problem over $R^{\prime}$ to $R$ (see for example [undefo, Proposition 2.1]), which explains the deep interest in the study of diophantine sets. In an attempt to contribute towards an existential definition of $\mathbb{Z}$ in $\mathbb{Q}$, Koenigsmann showed in [undefp] that $\mathbb{Z}$ is universally defined in $\mathbb{Q}$ (i.e. it is the complement of a diophantine set), by developing methods and results that were later generalized to prove analogous statements on more arbitrary global fields (see for example [undefaa], [undefi], [undeff], [undefx], [undefb]). Since the question about the undecidability of the positive existential theory of $\mathbb{Q}$ (or, more generally, a global field $K$) seems currently out of reach, research has tried an “intermediate step” approach, by studying intermediate subrings (some instances being [undefad], [undefj], [undefab], [undefg], [undefh]). Motivated by this approach, the author offered a $\forall\exists$-definition for every member of a sequence of sets that interpolates between a number field and its set of $S$-integers. These sets are known as the Campana points (see [undefaf]). Together with Handley, the author was able to extend his methods to give a similar approach for Darmon points in [undefag].
Recently Becher, Daans, and Dittmann were able to further generalize Koenigsmann’s methods to the case of algebraic function fields in one variable (see [undefa] and [undefc]). This motivated the author to transfer his previous results to this new scenario, and in this paper we prove:
@@
In the upcoming section we will see the specific definitions for Campana and Darmon points. It will be important to note from the beginning that these sets are given in terms of a finite set of $\mathbb{Z}$-valuations that are trivial on the base field. To be more specific, given an algebraic function field $F$ in one variable over a number field $K$, given $n\in\mathbb{Z}_{\geq 1}$ and given a finite set $S$ of $\mathbb{Z}$-valuations on $F$ that are trivial on $K$, we will define the sets $C_{F/K,S,n}$ (the set of $n$-Campana points of $F/K$ with respect to $S$) and $D_{F/K,S,n}$ (the set of $n$-Darmon points of $F/K$ with respect to $S$). As we will see, the first-order descriptions we will give will be uniform with respect to all possible choices of $S$ (cf. Remark 8.11 and Remark 8.13).
The above remark is important because, by [undefc, Corollary 5.9], if $K$ is a number field and $F$ is an algebraic function field in one variable over $K$, every computably enumerable relation on $F$ is $\exists\forall$-definable. In particular, since every decidable subset is both computably enumerable and has a computably enumerable complement, then all decidable subsets are both $\forall\exists$-definable and $\exists\forall$-definable. Since Campana points and Darmon points are decidable, we can get our first-order characterization a priori. The difference here is that our explicit approach towards a specific construction of their defining formulas will reveal to be uniform with respect to all finite subsets $S$ of $\mathbb{Z}$-valuations trivial on the base field in terms of which these sets are defined. More specifically, the sequence of Proposition 8.2, Proposition 8.3, and Proposition 8.6 will allow for a parametrization of all finite subsets of such places in a way that will allow for a uniform control on which of those sets $S$ is the one involved in the defining formula.
@@ 1.1. Understanding Campana and Darmon points
  • Darmon, if for all $\alpha\in\mathcal{A}$ with $\varepsilon_{\alpha}=1$ we have $n_{v}\left(\mathcal{D}_{\alpha},P\right)=0$, and for all $\alpha\in\mathcal{A}$ with $\varepsilon_{\alpha}<1$ we have $n_{\alpha}\mid n_{v}\left(\mathcal{D}_{\alpha},P\right)$.
For example, as shown in [undefag, §3.1], when $X=\mathbb{P}^{1}_{\mathbb{Q}}$, $\mathcal{X}=\mathbb{P}^{1}_{\mathbb{Z}}$, $D\coloneqq\left(1-\frac{1}{n}\right)\left\{x_{0}=0\right\}+\left\{x_{1}=0\right\}$ for some $n\in\mathbb{Z}_{\geq 1}$, and $S=\left\{\infty\right\}$, then the corresponding Darmon points in the affine line are the integral perfect $n$th powers up to sign. If instead $D\coloneqq\left(1-\frac{1}{n}\right)\left\{x_{0}=0\right\}+\left(1-\frac{1}{n}\right)\left\{x_{1}=0\right\}$ then we get rational $n$th powers up to sign.
If $K$ is any number field, $S$ is a finite set of places of $K$ containing the archimedean ones, $X=\mathbb{P}^{1}_{K}$, $\mathcal{X}=\mathbb{P}^{1}_{\mathcal{O}_{K,S}}$, $n\in\mathbb{Z}_{\geq 1}$ and $D=\left(1-\frac{1}{n}\right)\left\{x_{1}=0\right\}$, then calling $C_{K,S,n}$ and $D_{K,S,n}$ the corresponding Campana and Darmon points, respectively, as shown in [undefaf, §2.2] and [undefag, §3.2] we get
@@
These are the sets that were explored in [undefaf] and [undefag] and will be generalized in our new context.
@@ 1.1.1. Other contexts in which these sets have been studied
Even though these sets were mainly motivated by their interest in proving Manin-type conjectures and also in counting problems with height restrictions (see for instance [undef], [undefw], [undefac], [undefaj]), they have also been studied in the context of definability and decidability. For example, in [undefn] Garcia-Fritz, Pasten, and Pheidas showed that Kollár’s conjecture is incompatible with the diophantineness of Campana points in $\mathbb{C}\left(z\right)$, except for the trivial cases. We state the conjecture below. For each $n\in\mathbb{Z}_{\geq 1}$ we let $\mathbb{C}\left[z\right]_{n}$ be the $\left(n+1\right)$-dimensional $\mathbb{C}$-vector subspace of $\mathbb{C}\left[z\right]$ spanned by $\left\{1,x,\cdots,x^{n}\right\}$. We endow this set with the Zariski topology of $\mathbb{C}^{n+1}$.
@@
If $D$ is a diophantine subset of $\mathbb{C}\left(z\right)$ containing a non-empty Zariski open subset of $\mathbb{C}\left[z\right]_{n}$ for infinitely many $n\in\mathbb{Z}_{\geq 1}$, then $\mathbb{C}\left(z\right)\setminus D$ is finite.
@@
$C_{S,n}$ $\coloneqq\left\{0\right\}\cup\left\{f\in\mathbb{C}\left(z\right)^{\times}:\forall\alpha\in\mathbb{P}^{1}_{\mathbb{C}}\left(\mathbb{C}\right)\setminus S\,\left(\nu_{\alpha}\left(f\right)\in\mathbb{Z}_{\geq 0}\cup\mathbb{Z}_{\leq-n}\right)\right\},$
$D_{S,n}$ $\coloneqq\left\{0\right\}\cup\left\{f\in\mathbb{C}\left(z\right)^{\times}:\forall\alpha\in\mathbb{P}^{1}_{\mathbb{C}}\left(\mathbb{C}\right)\setminus S\,\left(\nu_{\alpha}\left(f\right)\in\mathbb{Z}_{\geq 0}\cup n\mathbb{Z}\right)\right\},$
@@
It is clear that $C_{S,1}=D_{S,1}=\mathbb{C}\left(z\right)$ is diophantine in $\mathbb{C}\left(z\right)$. Moreover, $C_{\emptyset,\infty}=D_{\emptyset,\infty}=\mathbb{C}$ is also diophantine in $\mathbb{C}\left(z\right)$ because, in general, if $L$ is a large field and $F$ is an algebraic function field in one variable over $L$, then $L$ is diophantine over $F$, by [undefq, Theorem 2]. We recall that a field $L$ is large if and only if every smooth curve over $L$ has either infinitely many or none $L$-rational points.
In [undefn, Proposition 1.13] Garcia-Fritz, Pasten, and Pheidas showed that, in the case of Campana points, these are the only instances of diophantineness that can occur if Conjecture 1.2 is true. For the sake of completeness, let us now show how this argument can be adapted to get an analogous proposition for Darmon points.
@@ 3. $\mathbb{Z}$-valuations on algebraic function fields in one variable
Recall that a $\mathbb{Z}$-valuation $v$ on a field $K$ is a discrete valuation such that for every $n\in\mathbb{Z}$ there exists $x\in K$ such that $n=v\left(x\right)$. If $L$ is a finite field extension of $K$ and $w$ is a $\mathbb{Z}$-valuation on $L$, we say that $w$ lies over $v$ (or that $v$ lies under $w$) if and only if $\mathcal{O}_{w}\cap K=\mathcal{O}_{v}$. In this case, there exists $e\left(w\mid v\right)\in\mathbb{Z}_{\geq 1}$ such that $w\left(x\right)=e\left(w\mid v\right)v\left(x\right)$ for every $x\in K$. The number $e\left(w\mid v\right)$ is called the ramification index of $w$ over $v$. We say that $v$ is unramified in $L$ if $e\left(w\mid v\right)=1$ for every $\mathbb{Z}$-valuation on $L$ lying over $K$. It is well-known that any $\mathbb{Z}$-valuation on $K$ lies under only finitely many (and at least one) $\mathbb{Z}$-valuations on $L$.
We refer the reader to [undefy] and [undefl] for more properties of valuations. In our case, we will be considering $\mathbb{Z}$-valuations on algebraic function fields in one variable.
@@
$C_{F/K,S,n}$ $\coloneqq\left\{0\right\}\cup\left\{x\in F^{\times}:v\left(x\right)\in\mathbb{Z}_{\geq 0}\cup\mathbb{Z}_{\leq-n}\text{ for all $v\in\mathcal{V}\left(F/K\right)\setminus S$}\right\},$
$D_{F/K,S,n}$ $\coloneqq\left\{0\right\}\cup\left\{x\in F^{\times}:v\left(x\right)\in\mathbb{Z}_{\geq 0}\cup n\mathbb{Z}\text{ for all $v\in\mathcal{V}\left(F/K\right)\setminus S$}\right\}.$
@@ 4. Quadratic Forms and Witt Equivalence
Assume fist $\operatorname{Char}\left(K\right)\neq 2$. By the diagonalization theorem for quadratic forms on characteristic different from $2$ ([undefk, Proposition 7.29]) and the fact that $q$ is non-singular, $q$ is isometric to the orthogonal product of quadratic forms $K\to K$ given by $x\mapsto ax^{2}$ for every $x\in K$, where $a\in K^{\times}$. Calling $\left\langle a\right\rangle$ such a quadratic form for a given $a\in K^{\times}$, the first conclusion follows from the fact that $\left\langle a\right\rangle\perp\left\langle-a\right\rangle$ is hyperbolic ([undefk, Example 7.26]). The second conclusion is a direct consequence of hyperbolic forms always being even-dimensional.
The proof when $\operatorname{Char}\left(K\right)=2$ is essentially the same; with the additional task of dealing with the more complicated version of the orthogonal factorization of quadratic forms on fields of characteristic $2$ ([undefk, Proposition 7.31]), since we have to also consider orthogonal factors $K\times K\to K$ of the form $\left(x,y\right)\mapsto ax^{2}+xy+by^{2}$ with $a,b\in K$. Calling $\left[a,b\right]$ such a quadratic form for given $a,b\in K$, [undefk, Example 7.23] gives $\left[a,b\right]\perp\left[a,-b\right]\cong\left[a,0\right]\perp\mathbb{H}\left(K\right)$, and [undefk, Example 7.5] gives $\left[a,0\right]\cong\mathbb{H}\left(K\right)$. $\blacksquare$